1) If a hypothesis is conducted using 𝐠= 0.05, for which of the following p-values would the null hypothesis be rejected?


2) For each pair of ? and p-value, indicate whether the null hypothesis would be rejected.


3) In a test of H₀: µ = 80 against H₁: µ<80, the sample data yielded the test statistic z = 1.63. What is


the p-value for the test?


4) In a test of H₀: µ = 80 against H₁: µ ≠ 80, a sample of observations had a mean and standard


deviation s = 4.1. What is the p-value for this test?


5) In a test of H₀: µ = 72 against H₁: µ ≠ 72, the sample data yielded the test statistic z = 1.63. Find the p-value for the test.

1
Expert's answer
2022-06-07T19:13:19-0400

1)

e) 0.001


2)

a) "p=0.10>0.05=\alpha," don't reject

b) "p=0.05=0.05=\alpha," don't reject

c) "p=0.001<0.01=\alpha," reject

d) "p=0.10>0.01=\alpha," don't reject

e) "p=0.01<0.10=\alpha," reject


3)

Left-tailed test "p=P(Z<1.63)=0.948449"


4)

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}"


The p-value for two-tailed, "df=n-1" degrees of freedom, "t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}" can be defined if "\bar{x}, \mu, s, n" are given.


5)

Two-tailed test "p=2P(Z>1.63)=0.103101"